OpenStudy (anonymous):

1.)What are the x- and y-intercepts of the line? 2x + 3y = –12 a. x-intercept –6, y-intercept –4 b. x-intercept –6, y-intercept 4 c. x-intercept –4, y-intercept 6 d. x-intercept 4, y-intercept –6 2.) Select the equation of the line that passes through the point (2, –1) and has slope 3 in point-slope form. a. (y + 1) = 3(x – 2) b. (y – 1) = 3(x + 2) c. (x + 2) = 3(y – 1) d. (x – 2) = 3(y + 1) 3.) What is the slope of the line that passes through the points (3, –1) and (–2, –5)? a. 5/4 b. -5/4 c. 4/5 d. -4/5

OpenStudy (anonymous):

I found it out! 1. is choice A; 2. is choice A; and 3. is choice C

OpenStudy (the_fizicx99):

Multiply the x or y by 0 and divide, 2x + 3y = -12 2(0) +3y=-12 /3 /3 y = -4 x = -6

OpenStudy (anonymous):

first a ans is a second a third c

OpenStudy (the_fizicx99):

@ashlashbigash @vvv it against CoCO to give out answer..

OpenStudy (the_fizicx99):

CoC *

OpenStudy (anonymous):

What is CoC?

OpenStudy (the_fizicx99):

Code of Conduct of this website

OpenStudy (anonymous):

but why then how to reply to a question

OpenStudy (anonymous):

Isn't that the whole point of this website?

OpenStudy (the_fizicx99):

Giving an explanation, not the answer.. and No. It's to guide learners.